# The cone conjecture for primitive symplectic varieties over a field of characteristic zero and an application

### v1

@@ 3.3. The absolute nef cone conjecture
The inclusion $\overline{\mathop{\rm BAmp}\nolimits(X)}\subseteq\mathop{\overline{\rm Mov}}\nolimits(X)$ is clear. To prove the converse inclusion, it suffices to show that every element in the interior of $\mathop{\rm Mov}\nolimits(X)$ is nef on a birational model of $X$. Take $\alpha\in\mathop{\rm Mov}\nolimits^{\circ}(X)$ and let $L^{\prime}\subset L$ be a finitely generated subfield and $X^{\prime}$ a projective primitive symplectic variety over $\overline{L^{\prime}}\subset L$ with $X^{\prime}_{L}=X$ and $\alpha\in\mathop{\rm Mov}\nolimits^{\circ}(X^{\prime})$. Let D be an $\mathbb{R}$-Cartier divisor on $X^{\prime}$ representing the class of $\alpha$. After fixing an embedding $\overline{L^{\prime}}\hookrightarrow\mathbb{C}$, we can consider the complex projective primitive symplectic variety $X^{\prime}_{\mathbb{C}}$. Both $X^{\prime}$ and $X^{\prime}_{\mathbb{C}}$ have $\mathbb{Q}$-factorial terminal singularities by Lemma 3.8. Recall that we have isometries
@@
induced by pullback of line bundles by Lemma 3.9. Consequently, since $\mathop{\rm Mov}\nolimits^{\circ}(X^{\prime}_{\mathbb{C}})\subset\mathop{\rm Pos}\nolimits(X^{\prime}_{\mathbb{C}})\subset\mathop{\rm Big}\nolimits(X^{\prime}_{\mathbb{C}})$ by [undefat, Lemma 4.6, Lemma 4.7], the same is true for $X^{\prime}$.
@@
We claim that only flips appear in any $K_{X^{\prime}}+D=D$ log-MMP on $X^{\prime}$. By contradiction, suppose that a divisorial contraction $f$ appears. The base change $f_{\mathbb{C}}$ remains a divisorial contraction since the relative Picard rank of $f_{\mathbb{C}}$ remains equal to $1$ by [undefaaa, Proposition 3.1] (note that although the proposition is stated for smooth varieties, the arguments in the proof hold in greater generality). The rest of the argument follows the proof of [undefat, Proposition 5.6]: there exists a $D_{\mathbb{C}}$-negative curve $R$ on $X^{\prime}_{\mathbb{C}}$ which can be contracted, and which covers a divisor $E\subset X^{\prime}_{\mathbb{C}}$. By [undefat, Corollary 2.14], the curve $R$ deforms along the its Hodge locus and at the general point $t$ of the latter, the Picard number $\rho(X^{\prime}_{\mathbb{C},t})$ is $1$. This means that the class of the deformation $R_{t}$ of $R$ is dual to a positive multiple of the class of the deformation $E_{t}$ of $E$, which means that there is $\lambda>0$ such that $R=\lambda E^{\vee}$. Then $0>R\cdot D=\lambda q_{X^{\prime}_{\mathbb{C}}}(E,D)\geq 0$, which leads to a contradiction.
@@
Running an MMP with scaling for a suitable ample divisor [undeff, Corollary 1.4.2], we obtain a birational model $\phi:X^{\prime}\dashrightarrow Y$ which is an isomorphism in codimension $1$ and for which $\phi_{*}D$ is nef. Hence $\alpha\in\phi_{L}^{*}\mathop{\rm Nef}\nolimits(Y_{L})$, and the proof is complete. ∎
Definition 3.35.
Let $L$ be an algebraically closed field of characteristic $0$ and $X$ a projective primitive symplectic variety over $L$ such that $b_{2}(X)\geq 5$ and $X$ has $\mathbb{Q}$-factorial and terminal singularities. Let $\Sigma\subset\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$ be the set
[…]
Here, $\mathop{\rm Nef}\nolimits(Y)^{\vee}$ is the dual cone of $\mathop{\rm Nef}\nolimits(Y)$ in $\mathop{\rm Pic}\nolimits(Y)_{\mathbb{R}}$ with respect to the BBF form. Clearly, $\Sigma$ is invariant under the action of $\mathop{\rm Bir}\nolimits(X)\rtimes W_{X}$ by definition.
When $L=\mathbb{C}$, we observe that $\Sigma$ is the set of all the classes of primitive wall divisors as defined in [undefat, Definition 7.1], see also [undefat, Lemma 3.7] to understand why these classes are the only classes of wall divisors.
@@
Let $L\subset M$ be a field extension of algebraically closed field of characteristic $0$ and $X$ a projective primitive symplectic variety over $L$ with $\mathbb{Q}$-factorial terminal singularities. Denote by $p:X_{M}\to X$ the projection. The isomorphism $p^{*}:\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}\xrightarrow{\simeq}\mathop{\rm Pic}\nolimits(X_{M})_{\mathbb{R}}$ from Lemma 3.9 induces a bijection between $\mathop{\rm BAmp}\nolimits(X)$ and $\mathop{\rm BAmp}\nolimits(X_{M})$.
@@
First, note that $p^{*}\mathop{\rm BAmp}\nolimits(X)\subseteq\mathop{\rm BAmp}\nolimits(X_{M})$. Next, by Lemma 3.3, it is clear that $p^{*}\mathop{\overline{\rm Mov}}\nolimits(X)=\mathop{\overline{\rm Mov}}\nolimits(X_{M})$. Therefore, Lemma 3.34 implies that $p^{*}\mathop{\rm BAmp}\nolimits(X)$ is dense in $\mathop{\rm Mov}\nolimits(X_{M})$. Let $f:X_{L}\dashrightarrow Y$ be a birational map as in Definition 3.31 and consider the open convex subcone $f^{*}\mathop{\rm Amp}\nolimits(Y)$ of $\mathop{\rm Mov}\nolimits(X_{M})$. By density, there is a birational model $g:X\dashrightarrow Z$ as in Definition 3.31 such that $p^{*}g^{*}\mathop{\rm Amp}\nolimits(Z)\cap f^{*}\mathop{\rm Amp}\nolimits(Y)\neq\emptyset$. But now we notice that $p^{*}g^{*}\mathop{\rm Amp}\nolimits(Z)=g_{L}^{*}\mathop{\rm Amp}\nolimits(Z_{M})$ with $g_{M}:X_{M}\dashrightarrow Z_{M}$ the base change of $g$, hence
@@
Let $L$ be an algebraically closed field of characteristic $0$ and $X$ a projective primitive symplectic variety over $L$ with $b_{2}(X)\geq 5$ and $\mathbb{Q}$-factorial terminal singularities. Then, up to $L$-isomorphisms, $X$ has a finite number of $\mathbb{Q}$-factorial terminal birational models.
@@
Take a finitely generated subfield $L^{\prime}\subset L$ and let $X^{\prime}$ be a projective primitive symplectic variety over $\overline{L^{\prime}}\subset L$ such that $X^{\prime}_{L}=X$. We fix an embedding $\overline{L^{\prime}}\hookrightarrow\mathbb{C}$ and consider the projective primitive symplectic variety $X^{\prime}_{\mathbb{C}}$ with $\mathbb{Q}$-factorial terminal singularities and $b_{2}(X^{\prime}_{\mathbb{C}})\geq 5$. Since $X^{\prime}_{\mathbb{C}}$ has, up to $\mathbb{C}$-isomorphism, finitely many $\mathbb{Q}$-factorial and terminal birational models by [undefat, Corollary 1.4], $\mathop{\rm BAmp}\nolimits(X_{\mathbb{C}})$ is a finite union of pullbacks of ample cones. Furthermore, these cones are precisely the connected components of $\mathop{\rm BAmp}\nolimits(X^{\prime}_{\mathbb{C}})$. Using Lemma 3.36, we deduce that the number of connected components of $\mathop{\rm BAmp}\nolimits(X^{\prime})$, and therefore of $\mathop{\rm BAmp}\nolimits(X)$, is finite. This shows that the union defining the birational ample cone of $X$ is in fact finite, and we can therefore deduce that the number of models is finite. ∎
Thanks to this information, we deduce as in the complex case that the square with respect to the BBF form of the primitive wall divisor classes is uniformly bounded from below.
@@
Let $L$ be an algebraically closed field of characteristic $0$ and $X$ a projective primitive symplectic variety over $L$ such that $b_{2}(X)\geq 5$ and $X$ has $\mathbb{Q}$-factorial and terminal singularities. Then there exists $B>0$ such that $q_{X}(\alpha)\geq-B$ for all $\alpha$ in $\Sigma$.
@@
Given that for $\alpha=\omega\cdot f^{*}(e)\in\Sigma$, we have $q_{X}(\alpha)=q_{X}(f^{*}(e))$, it suffices to prove the statement for elements of the form $f^{*}(e)\in\Sigma$. Let $n\geq 1$ be the number of $\mathbb{Q}$-factorial and terminal models of $X$ up to $L$-isomorphisms. For all $i\in\llbracket 1,n\rrbracket$, let us fix a representative $f_{i}:X\dashrightarrow Y_{i}$. Take a finitely generated subfield $L^{\prime}\subset L$ such that the $f_{i}$, the $Y_{i}$ and $X$ are defined on $\overline{L^{\prime}}\subset L$. We fix an embedding $\overline{L^{\prime}}\hookrightarrow\mathbb{C}$ and consider the projective primitive symplectic variety $X_{\mathbb{C}}$. On $X_{\mathbb{C}}$, there is a $B>0$ such that the statement of the proposition is true thanks to [undefat, Proposition 7.7]. Each $f^{*}(e)\in\Sigma$ is equal to $f_{i}^{*}(e_{i})$ for a certain index $i\in\llbracket 1,n\rrbracket$, with $e_{i}$ as in Definition 3.35. By base change, $f^{*}(e)$ corresponds in $X_{\mathbb{C}}$ to a class of a wall divisor and therefore $q_{X_{\mathbb{C}}}(f^{*}(e))\geq-B$ (it should be understood here that $f^{*}(e)$ is seen as an element of $\mathop{\rm Pic}\nolimits(X_{\mathbb{C}})_{\mathbb{R}}$ thanks to the various base changes). But now the projections associated with the base changes induce isometries with respect to the BBF forms by Lemma 3.9, which completes the proof. ∎
We can now prove Theorem A.(2) by adapting the proof of [undefat, Theorem 1.2].
@@
Let $\Pi\subset$ be a rational polyhedral cone which is a fundamental domain for the action of $\mathop{\rm Bir}\nolimits(X)$ on $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$ as in Theorem A.(1). By Proposition 3.38 and [undefay, Proposition 3.4], the set
@@
As $\mathop{\rm Nef}\nolimits^{+}(X)$ is a union of $\mathop{\rm Aut}\nolimits(X)$-translates of the finitely many rational polyhedral cones $\Pi_{i}$ for $i\in I_{X}$, we deduce that ${\mathcal{D}}_{y}$ is a fundamental rational polyhedral domain for the action of $\mathop{\rm Aut}\nolimits(X)$ on $\mathop{\rm Nef}\nolimits^{+}(X)$ by [undefat, Lemma 7.8], and this completes the proof. ∎

### v2

@@ 3.3. The absolute nef cone conjecture
The inclusion $\overline{\mathop{\rm BAmp}\nolimits(X)}\subseteq\mathop{\overline{\rm Mov}}\nolimits(X)$ is clear. To prove the converse inclusion, it suffices to show that every rational element in the interior of $\mathop{\rm Mov}\nolimits(X)$ is nef on a birational model of $X$. Take $\alpha\in\mathop{\rm Mov}\nolimits^{\circ}(X)\cap\mathop{\rm Pic}\nolimits(X)_{\mathbb{Q}}$, and let $D$ be a $\mathbb{Q}$-Cartier divisor on $X$ representing the class of $\alpha$. Let $L^{\prime}\subset L$ be a finitely generated subfield, $X^{\prime}$ a projective primitive symplectic variety over $\overline{L^{\prime}}\subset L$ and $D^{\prime}$ a $\mathbb{Q}$-Cartier divisor on $X^{\prime}$ such that $X=X^{\prime}_{L}$ and $D=D^{\prime}_{L}$. Fix an embedding $\overline{L^{\prime}}\hookrightarrow\mathbb{C}$, and consider the complex projective primitive symplectic variety $X^{\prime}_{\mathbb{C}}$. Both $X^{\prime}$ and $X^{\prime}_{\mathbb{C}}$ have $\mathbb{Q}$-factorial terminal singularities by Lemma 3.8. Recall that we have isometries
$\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits(X^{\prime})_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits(X^{\prime}_{\mathbb{C}})_{\mathbb{R}}$ (3.7)
@@
induced by pullback of line bundles by Lemma 3.9. The movable cones of $X$, $X^{\prime}$ and $X^{\prime}_{\mathbb{C}}$ correspond bijectively through the above isometries by Lemma 3.3, therefore $D^{\prime}$ and $D^{\prime}_{\mathbb{C}}$ lie respectively in $\mathop{\rm Mov}\nolimits^{\circ}(X^{\prime})\cap\mathop{\rm Pic}\nolimits(X^{\prime})_{\mathbb{Q}}$ and $\mathop{\rm Mov}\nolimits^{\circ}(X^{\prime}_{\mathbb{C}})\cap\mathop{\rm Pic}\nolimits(X^{\prime}_{\mathbb{C}})_{\mathbb{Q}}$. Since $\mathop{\rm Mov}\nolimits^{\circ}(X^{\prime}_{\mathbb{C}})\subset\mathop{\rm Pos}\nolimits(X^{\prime}_{\mathbb{C}})\subset\mathop{\rm Big}\nolimits(X^{\prime}_{\mathbb{C}})$ by [undefas, Lemma 4.6, Lemma 4.7], the same is true for $X^{\prime}$ once again thanks to (3.7) and Lemma 3.3, thus $D^{\prime}$ is big.
@@
We claim that only flips appear in any $K_{X^{\prime}}+D^{\prime}=D^{\prime}$ log-MMP on $X^{\prime}$. Suppose by contradiction that a divisorial contraction $f$ appears. The base change $f_{\mathbb{C}}$ remains a divisorial contraction since the relative Picard rank of $f_{\mathbb{C}}$ remains equal to one by [undefaaa, Proposition 3.1] (note that although the proposition is stated for smooth varieties, the arguments in the proof hold in greater generality). The rest of the argument follows the proof of [undefas, Proposition 5.6]: there exists a $D^{\prime}_{\mathbb{C}}$-negative curve $R$ on $X^{\prime}_{\mathbb{C}}$ which can be contracted, and which covers a divisor $E\subset X^{\prime}_{\mathbb{C}}$. By [undefas, Corollary 2.14], the curve $R$ deforms along the its Hodge locus and at the general point $t$ of the latter, the Picard number $\rho(X^{\prime}_{\mathbb{C},t})$ is one. This means that the class of the deformation $R_{t}$ of $R$ is dual to a positive multiple of the class of the deformation $E_{t}$ of $E$, which means that there is $\lambda>0$ such that $R=\lambda E^{\vee}$. Then $0>R\cdot D^{\prime}_{\mathbb{C}}=\lambda q_{X^{\prime}_{\mathbb{C}}}(E,D^{\prime}_{\mathbb{C}})\geq 0$, which leads to a contradiction.
@@
Running an MMP with scaling for a suitable ample divisor [undefav, Corollary 21.9], we obtain a birational model $\phi:X^{\prime}\dashrightarrow Y$ for which $\phi_{*}D^{\prime}$ is nef, and that completes the proof. ∎
@@
Let $L\subset M$ be a field extension of algebraically closed field of characteristic $0$ and $X$ a projective primitive symplectic variety over $L$ with $\mathbb{Q}$-factorial terminal singularities. Denote by $p:X_{M}\to X$ the projection. The isomorphism $p^{*}:\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}\to\mathop{\rm Pic}\nolimits(X_{M})_{\mathbb{R}}$ from Lemma 3.9 induces a bijection between $\mathop{\rm BAmp}\nolimits(X)$ and $\mathop{\rm BAmp}\nolimits(X_{M})$.
@@
By base change, it is clear that $p^{*}\mathop{\rm BAmp}\nolimits(X)\subseteq\mathop{\rm BAmp}\nolimits(X_{M})$. For the reverse inclusion, observe that $p^{*}\mathop{\overline{\rm Mov}}\nolimits(X)=\mathop{\overline{\rm Mov}}\nolimits(X_{M})$ by Lemma 3.3. Therefore, Lemma 3.35 implies that $p^{*}\mathop{\rm BAmp}\nolimits(X)$ is dense in $\mathop{\rm Mov}\nolimits(X_{M})$. Let $f:X_{M}\dashrightarrow Y$ be a birational map as in Definition 3.32. Since $f^{*}\mathop{\rm Amp}\nolimits(Y)$ is an open subset of $\mathop{\rm Mov}\nolimits(X_{M})$, there exists a birational model $g:X\dashrightarrow Z$ as in Definition 3.32 such that
$p^{*}g^{*}\mathop{\rm Amp}\nolimits(Z)\cap f^{*}\mathop{\rm Amp}\nolimits(Y)\neq\emptyset.$
If $g_{M}:X_{M}\dashrightarrow Z_{M}$ denotes the base change of $g$, we have $p^{*}g^{*}\mathop{\rm Amp}\nolimits(Z)=g_{M}^{*}\mathop{\rm Amp}\nolimits(Z_{M})$, therefore
@@
Let $L$ be an algebraically closed field of characteristic zero and $X$ a projective primitive symplectic variety over $L$ with $b_{2}(X)\geq 5$ and $\mathbb{Q}$-factorial terminal singularities. Then, $X$ has a finite number of marked minimal models $X^{\prime}$ (Definition 4.8) over $L$.
@@
Take a finitely generated subfield $L^{\prime}\subset L$ and let $X^{\prime}$ be a projective primitive symplectic variety over $\overline{L^{\prime}}\subset L$ such that $X=X^{\prime}_{L}$. Fix an embedding $\overline{L^{\prime}}\hookrightarrow\mathbb{C}$, and consider the complex projective primitive symplectic variety $X^{\prime}_{\mathbb{C}}$. By [undefas, Corollary 1.4], $X^{\prime}_{\mathbb{C}}$ has finitely many marked minimal models, therefore $\mathop{\rm BAmp}\nolimits(X^{\prime}_{\mathbb{C}})$ is a finite union of pullbacks of ample cones. Moreover, these cones are precisely the connected components of $\mathop{\rm BAmp}\nolimits(X^{\prime}_{\mathbb{C}})$. Using Lemma 3.36, we deduce that $\mathop{\rm BAmp}\nolimits(X^{\prime})$ and $\mathop{\rm BAmp}\nolimits(X)$ have finitely many connected components. This shows that the union defining the birational ample cone of $X$ is in fact finite, and it follows that the number of marked minimal models of $X$ is finite. ∎
Definition 3.38.
Let $L$ be an algebraically closed field of characteristic $0$, and let $X$ be a projective primitive symplectic variety over $L$ with $b_{2}(X)\geq 5$ and $\mathbb{Q}$-factorial terminal singularities. Let $\Sigma(X)\subset\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}$ be the set
[…]
Here, $\mathop{\rm Nef}\nolimits(Y)^{\vee}$ is the dual cone of $\mathop{\rm Nef}\nolimits(Y)$ in $\mathop{\rm Pic}\nolimits(Y)_{\mathbb{R}}$ with respect to the Beauville-Bogomolov-Fujiki form.
As in the complex case, the square with respect to the Beauville-Bogomolov-Fujiki form of the elements of $\Sigma(X)$ is uniformly bounded from below.
@@
Let $L$ be an algebraically closed field of characteristic $0$, and let $X$ be a projective primitive symplectic variety over $L$ with $b_{2}(X)\geq 5$ and $\mathbb{Q}$-factorial terminal singularities. Then there exists $B>0$ such that $q_{X}(x)\geq-B$ for all $x$ in $\Sigma(X)$.
@@
By Corollary 3.37, $X$ has a finite number $n\geq 1$ of marked minimal models. For each $i\in\{1,\dotsc,n\}$, let $f_{i}:X\dashrightarrow Y_{i}$ be a birational map defining a marking. Take a finitely generated subfield $L^{\prime}\subset L$ such that $X$, the $Y_{i}$ and the $f_{i}$ are defined over $\overline{L^{\prime}}\subset L$. This means that there exist projective primitive symplectic varieties $X^{\prime}$ and $Y^{\prime}_{i}$ over $\overline{L^{\prime}}\subset L$, together with birational maps $f^{\prime}_{i}:X^{\prime}\dashrightarrow Y^{\prime}_{i}$, such that $X=X^{\prime}_{L}$, $Y_{i}=Y^{\prime}_{i,L}$ and $f_{i}=f^{\prime}_{i,L}$. By Lemma 3.8, $X^{\prime}$ and the $Y^{\prime}_{i}$ are projective primitive symplectic varieties with $\mathbb{Q}$-factorial terminal singularities. Fix an embedding $\overline{L^{\prime}}\hookrightarrow\mathbb{C}$ and consider the base changes $X^{\prime}_{\mathbb{C}}$, $Y^{\prime}_{i,\mathbb{C}}$ and $f^{\prime}_{i,\mathbb{C}}$. For the same reason, $X^{\prime}_{\mathbb{C}}$ and the $Y^{\prime}_{i,\mathbb{C}}$ also have $\mathbb{Q}$-factorial terminal singularities. By Lemma 3.9, there are isometries
$\mathop{\rm Pic}\nolimits(X)_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits(X^{\prime})_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits(X^{\prime}_{\mathbb{C}})_{\mathbb{R}},\penalty 10000\ \text{and}\penalty 10000\ \mathop{\rm Pic}\nolimits(Y_{i})_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits(Y^{\prime}_{i})_{\mathbb{R}}\simeq\mathop{\rm Pic}\nolimits(Y^{\prime}_{i,\mathbb{C}})_{\mathbb{R}}$
for all $i\in\{1,\dotsc,n\}$. For each index $i$, the nef cones of $Y_{i}$, $Y^{\prime}_{i}$ and $Y^{\prime}_{i,\mathbb{C}}$ correspond bijectively under these isometries by Lemma 3.3. Moreover, since isomorphisms descend along base changes induced by field extensions by [undefaau, Tag 02L4], two distinct marked minimal models of $X^{\prime}$ cannot become isomorphic after base change to $L$ or to $\mathbb{C}$. Therefore, the projection $X\to X^{\prime}$ induces a bijection $\Sigma(X)\simeq\Sigma(X^{\prime})$, and the projection $X^{\prime}_{\mathbb{C}}\to X^{\prime}$ induces an injective map $\Sigma(X^{\prime})\hookrightarrow\Sigma(X^{\prime}_{\mathbb{C}})$ such that the composite
$\Sigma(X)\hookrightarrow\Sigma(X^{\prime}_{\mathbb{C}})$
preserves the Beauville-Bogomolov-Fujiki form. By [undefas, Proposition 7.7], there exists a real number $B>0$ such that, for every $x\in\Sigma(X^{\prime}_{\mathbb{C}})$, we have $q_{X^{\prime}_{\mathbb{C}}}(x)\geq-B$. In particular, this inequality holds for every element of $\Sigma(X)$, which completes the proof. ∎
We can now prove Theorem A.(2) by adapting the proof of [undefas, Theorem 1.2].
@@
Let $\Pi\subset$ be a rational polyhedral cone which is a fundamental domain under the action of $\mathop{\rm Bir}\nolimits(X)$ on $\mathop{\overline{\rm Mov}}\nolimits^{+}(X)$ as in Theorem A.(1). By Proposition 3.39 and [undefay, Proposition 3.4], the set
@@
As $\mathop{\rm Nef}\nolimits^{+}(X)$ is a union of $\mathop{\rm Aut}\nolimits(X)$-translates of the finitely many rational polyhedral cones $\Pi_{i}$ for $i\in I_{X}$, we deduce that ${\mathcal{D}}_{y}$ is a fundamental rational polyhedral domain under the action of $\mathop{\rm Aut}\nolimits(X)$ on $\mathop{\rm Nef}\nolimits^{+}(X)$ by [undefas, Lemma 7.8], and this completes the proof. ∎